Suppose three coins are tossed find each probability. Please do all
1.)P( exactly 1 tail)
2.)P(exactly 2 heads)
3.)P(3 tails)
4.)P(all the same)
5.)P(1 or more tails)
6.)P(3 heads)

Please help me thanks!!!
P.S Please do all



Answer by stanbon(48568)

Suppose three coins are tossed find each probability. Please do all
1.)P( exactly 1 tail) = 3C1(1/2)^1(1/2)^2 = 3*(1/8) = 3/8
----------------------------
2.)P(exactly 2 heads) = 3C2(1/2)^2(1/2) = 3*(1/8) = 3/8
-------
Note: exactly 1 tail means the same as exactly 2 heads
----------------------------
3.)P(3 tails) = 3C3(1/2)^3 = 1*(1/8) = 1/8
--------------------------
4.)P(all the same) 2*(1/8) = 1/4
-----------
5.)P(1 or more tails) = 1 - P(no tails) = 1-(1/8) = 7/8
------------------
6.)P(3 heads) = (1/8)
=======================
Cheers,
Stan H.